# LaV-CoT: Language-Aware Visual CoT with Multi-Aspect Reward Optimization for Real-World Multilingual VQA

### v3

@@ 7. Online A/B Test
To evaluate the real-world effectiveness of our proposed approach, we integrated LaV-CoT into a full-scale intelligent document processing system that provides multilingual visual question answering capabilities. The system operates in realistic enterprise environments, handling diverse document types, complex layouts, and multiple languages. Unlike traditional pipelines, which often rely on heuristic rules or single-stage reasoning models, our method employs structured multi-stage reasoning to sequentially process document content, summarize text segments, identify languages, generate object-level image captions, and perform step-by-step reasoning to answer complex queries.
We conducted a online A/B test comparing LaV-CoT with the existing production pipeline. Key business and user-centric metrics were tracked, including answer acceptance rate and user satisfaction scores collected through post-interaction feedback. The results demonstrate that LaV-CoT substantially outperforms the baseline: the answer acceptance rate increased by 8.7%, and the user satisfaction score improved by 12.4%, confirming that structured multi-stage reasoning enhances both accuracy and overall user experience in practical document understanding scenarios.
@@ 8. Conclusion
In this paper, we introduced LaV-CoT, a novel framework for multilingual multimodal question answering in real-world scenarios. By integrating language-aware visual Chain-of-Thought reasoning with structured supervision and a novel GRPO-based multi-aspect reward design, LaV-CoT effectively addresses the limitations of existing approaches that struggle with multilingual alignment, language inconsistency, and complex reasoning tasks. Our reward formulation jointly evaluates language consistency, structural accuracy, and semantic alignment, enabling more interpretable reasoning and significantly improving both model stability and generalization. Extensive experiments across diverse benchmarks demonstrate that LaV-CoT not only outperforms strong open-source baselines of similar scale, but also rivals much larger models and closed-source frontier systems. Furthermore, its deployment on a large-scale Intelligent Document Processing platform confirms its practical value, achieving notable improvements in accuracy, user satisfaction, and cost-effectiveness. Looking ahead, we plan to extend LaV-CoT to broader low-resource languages and domain-specific applications, while further exploring advanced reward modeling to enhance the robustness and inclusiveness of multilingual multimodal reasoning systems.
9. Limitations
Despite the promising results, LaV-CoT still has several limitations:
  • Sensitivity to multilingual input quality. The effectiveness of LaV-CoT is contingent on the quality of the multimodal inputs. When document images contain heavy language mixing across different scripts, the reasoning pipeline may struggle to maintain language consistency and semantic fidelity.
  • Coverage of low-resource languages. Our current training primarily relies on open-source multilingual VQA datasets that focus on high- and medium-resource languages. However, high-quality datasets for truly low-resource or less common languages remain scarce, and their collection and construction are necessary for broader inclusiveness.
  • Limited exploration of fast-slow reasoning integration. LaV-CoT is currently designed with a slow, multi-step reasoning pipeline to enhance interpretability. While effective, it does not yet incorporate fast thinking strategies or hybrid fast-slow reasoning mechanisms, which we plan to explore in future work to further improve efficiency and adaptability.
@@ 10. Ethical Use of Data
Acknowledgements.
This work was supported by ZoloZ Pte. Ltd.
Appendix A Dataset Statistics
Table 3 provides an overview of the datasets used in our experiments. These benchmarks span multiple languages and cover a diverse range of tasks, ensuring a comprehensive evaluation of multilingual visual question answering.
Dataset Language Size
COCO2017 EN,ZH,PT,AR,TR,RU 73k
Visual-Genome EN,ZH,PT,AR,TR,RU 18k
GQA EN,ZH,PT,AR,TR,RU 15k
OCR-VQA EN,ZH,PT,AR,TR,RU 16k
TextVQA EN,ZH,PT,AR,TR,RU 4.7k
Llava-Pretrain EN,ZH,PT,AR,TR,RU 0.6k
MTVQA AR,DE,FR,IT,JA,KO,RU,TH,VI 21k
Appendix B Prompt Design
B.1. Generator Prompt
Figure 6. Instructions for generating Vanilla CoT.:
Figure 7. Instructions for correcting error step.:
B.2. Evaluator Prompt
Figure 8. Instructions for evaluating cot step.:
B.3. Inference Prompt
Figure 9. Instructions for LaV-CoT inference on open-ended questions.:
Figure 10. Instructions for LaV-CoT inference on MCQ.:
@@ Appendix C More Real-world Cases
Figure 11. Arabic Visa demo case.:
Figure 12. Indonesian ID demo case.:

### v4

@@ 7. Online A/B Test
We conducted an online A/B test comparing LaV-CoT with the existing production pipeline on our Intelligent Document Processing platform  [^1] https://www.zoloz.com/realdoc for two weeks. Key business and user-centric metrics were tracked, including answer acceptance rate and user satisfaction scores collected through post-interaction feedback. The results demonstrate that LaV-CoT substantially outperforms the baseline: the answer acceptance rate increased by 8.7%, and the user satisfaction score improved by 12.4%, confirming that structured multi-stage reasoning enhances both accuracy and overall user experience in practical document understanding scenarios.
Figure 5. Ablation study of multi-aspect reward configurations on MTVQA. Here R_Lang represents Language Consistency Reward, R_Count represents Text Segments and Object Count Reward, R_Answer represents Edit Distance of Final Answer Reward, and R_Format represents Format Reward.
ablation rewards
@@ 8. Conclusion
In this paper, we present LaV-CoT, a novel framework for multilingual Text-Centric VQA. By combining language-aware visual Chain-of-Thought reasoning with structured supervision and GRPO-based multi-aspect rewards, LaV-CoT improves interpretability, visual-text alignment, and language consistency. Extensive experiments across diverse benchmarks show that LaV-CoT outperforms comparable open-source models and rivals larger or proprietary ones. Furthermore, deployment on an Intelligent Document Processing platform demonstrates practical gains in accuracy, user satisfaction, and efficiency.
While promising, our approach has several limitations. Coverage of low-resource languages remains constrained by the scarcity of high-quality multilingual data. Additionally, the current framework focuses on slow, step-by-step reasoning and does not yet support fast or hybrid reasoning. In future work, we aim to extend LaV-CoT to broader low-resource and domain-specific scenarios, incorporate fast and hybrid reasoning mechanisms, and enhance multi-aspect reward modeling to build more robust and inclusive multimodal reasoning systems.
@@ 9. Ethical Use of Data
10. Acknowledgments
This work was supported by the Ant Group Postdoctoral Programme. This work was also supported by the Japan Science and Technology Agency (JST) and the Agency for Science, Technology and Research (A*STAR) under the Japan-Singapore Joint Call (Project No. R24I6IR133).
Appendix A Prompt Design
A.1. Generator Prompt
Figure 6. Instructions for generating Vanilla CoT.
Vanilla CoT Prompt.
Figure 7. Instructions for correcting error steps.
Correcting error CoT Prompt.
A.2. Evaluator Prompt
Figure 8. Instructions for evaluating CoT steps.
Evaluating CoT Prompt.
A.3. Inference Prompt
Figure 9. Instructions for LaV-CoT inference on open-ended questions.
Instructions for LaV-CoT inference on open-ended questions.
Figure 10. Instructions for LaV-CoT inference on MCQ.
Instructions for LaV-CoT inference on MCQ.
Appendix B Additional Performance Comparison
Model DocVQA ChartQA OCRBenchv2(en/cn)
Qwen2.5-VL-3B 93.9 83.9 52.1/54.3
Qwen2.5-VL-7B 94.1 84.2 53.2/56.1
GPT-4o-2024-05-13 94.4 85.1 55.8/57.6
Gemini-2.5-flash 94.9 86.0 56.0/57.9
LaV-CoT-3B(SFT) 94.5 85.4 55.2/57.6
LaV-CoT-3B(SFT+GRPO) 95.0 86.3 56.4/58.2
Table 4. Comparison of different models on standard document understanding benchmarks.
@@ Appendix C More Real-world Cases
Figure 11. Arabic Visa demo case.
Arabic visa demo case.
Figure 12. Indonesian ID demo case.
Indonesian ID demo case.